The reaction between solid white phosphorus and oxygen produces solid

tetraphosphorusdecoxide (P4O10).

a. Write the balanced chemical equation of the reaction. ( 5 points)

b. Determine the mass of tetraphosphorusdecoxide (P4O10) formed if 25.0 g of

phosphorus (P4) and 50.0 g of oxygen (O2) are combined. ( 5 points)

c. How much of the excess reactant remains after the reaction stops? (5 points)


1
Expert's answer
2022-04-29T14:41:04-0400

P4+5O2=P4O10

m(P4O10)=0.2*284=56.8 g

c)m(O2)=50-18=32 g